OpenStudy (anonymous):

Martha cycled 14 miles each day on Saturday and Monday and 13 miles each day on Tuesday and Thursday. How many miles did she cycle in all?

OpenStudy (confluxepic):

14 on Sunday 14 on Monday 13 on Tuesday 13 on Thursday Add them all up 14+14+13+13 =28+26= 54 miles in all.